# A Smooth Approximation Framework for Weakly Convex Optimization

### v1

@@ 1 Introduction
Here, $f$ is a nonsmooth nonconvex function, and $r$ is a convex, lower-bounded, and lower semi-continuous function. We assume that $f$ is $\rho$-weakly convex for some $\rho>0$; that is, the function $f(\mathbf{x})+\frac{\rho}{2}\|\mathbf{x}\|^{2}$ is convex. We also assume that $r$ is prox-friendly, meaning its associated proximal operator can be computed efficiently. Weakly convex optimization problems frequently arise in various data-driven applications, including signal processing [duchi2019solving], machine learning [charisopoulos2021low], and reinforcement learning [wang2023policy].
The nonsmoothness of $f$ prevents the direct application of classical complexity results for smooth nonconvex optimization [ghadimi2016mini, saeed-lan-nonconvex-2013]. Early progress in weakly convex optimization was achieved using double-loop proximal algorithms that require solving auxiliary subproblems at each iteration [damek19Proximally, drusvyatskiy2019efficiency]. A major breakthrough came from davis2019stochastic, who proved that single-loop methods, such as the (stochastic) subgradient method, achieve an $\mathcal{O}(1/\varepsilon^{4})$ complexity bound for finding an approximate stationary point. Their work reveals an implicit smoothing effect in nonsmooth optimization: algorithms like the proximal subgradient method can be interpreted as optimizing the Moreau envelope $\phi^{\hat{\rho}}(\mathbf{x})$ ($\hat{\rho}>\rho$), a smooth approximation of $\phi$. Notably, $\phi$ and its Moreau envelope $\phi^{\hat{\rho}}$ share the same stationary points, and the gradient of the envelope provides a natural measure of approximate stationarity for (1). This insight has established the Moreau envelope as a central tool in the algorithm analysis for weakly convex problems, influencing a wide range of subsequent work [mai2020convergence, deng2021minibatch, gao24-stochastic, chen2021distributed, li2024revisiting, zhu2023unified].
@@
Explicit smoothing has a long history [chen2012smoothing]. In particular, Nesterov’s pioneering work [nesterov2005smooth] showed that applying accelerated methods to a suitably smoothed approximation of a structured convex problem improves the complexity of finding an $\varepsilon$-optimal solution from $\mathcal{O}(1/\varepsilon^{2})$ to $\mathcal{O}(1/\varepsilon)$. Building on this idea, Beck2012smoothing developed a unified smoothing framework that integrates multiple smoothing schemes with accelerated gradient methods. However, both [nesterov2005smooth] and [Beck2012smoothing] restrict their analyses to convex problems. While earlier works have examined the asymptotic behavior of smoothing in nonconvex settings [chen2012smoothing, xu2009smooth], a rigorous complexity theory for explicit smoothing in the weakly convex regime has remained underdeveloped.
@@ 1.1 Contributions
To address this gap, we extend the smooth approximation framework of Beck2012smoothing from the convex setting to the broader class of weakly convex problems. In particular, our framework seamlessly incorporates two prominent smoothing techniques as special cases: 1) Moreau envelope smoothing, and 2) Nesterov-type smoothing for compositions of convex functions with smooth mappings. A key advantage of this generalization arises in stochastic optimization, where one can smooth the function associated with each stochastic sample individually and then perform minibatching. This strategy can provide computational benefits over model-based minibatching algorithms [davis2019stochastic, deng2021minibatch], which require solving a potentially expensive proximal subproblem at every iteration. Another important consequence is that Moreau envelope smoothing becomes substantially more practical: it can now be used not only for convergence analysis [davis2019stochastic], but also as an effective and scalable algorithmic tool.
It is important to emphasize that the smooth approximation we propose, akin to that of Beck2012smoothing, describes a broad class of surrogate functions and is not tied to any particular smoothing technique. We establish quantitative relationships connecting the stationarity measures of the smoothed problem (2) to those of the original nonsmooth problem (1). A key implication is that the design of the smooth surrogate can be decoupled from the design of the optimization algorithm. This stands in contrast to many prior smoothing strategies for weakly convex optimization, which often rely on exploiting structural properties of the underlying problem [bohm2021variable, drusvyatskiy2019efficiency].
@@
A further contribution of our work is the relaxation of the global Lipschitz smoothness assumption [Beck2012smoothing] in the design of smooth approximation functions. As a result, we can handle objectives that are not globally Lipschitz, thereby substantially broadening the class of problems to which our framework applies. For example, in real phase retrieval, the loss $\ell(\mathbf{x})=|\langle\mathbf{a},\mathbf{x}\rangle|^{2}-b$ has an unbounded subgradient, and standard Nesterov-type smoothing produces approximations that fail to satisfy global Lipschitz smoothness. To overcome this limitation, we introduce a line search-based accelerated gradient method for convex optimization problems without global Lipschitz smoothness. The use of line search ensures that accelerated convergence rates for convex problems remain attainable even when global smoothness is absent. Under mild assumptions, we show that the inexact proximal point method, equipped with this accelerated solver for its subproblems, still achieves an $\mathcal{O}(1/\varepsilon^{3})$ complexity guarantee. This result markedly improves upon the previous $\mathcal{O}(1/\varepsilon^{4})$ complexity achieved by non-Lipschitz subgradient methods [grimmer2019convergence, zhu2023unified]. Finally, our theoretical findings are supported by numerical experiments. Our results demonstrate that smoothing leads to smoother convergence behavior and outperforms standard subgradient methods on robust nonlinear regression problems.
@@ 1.2 Related works
The $\mathcal{O}(1/\varepsilon^{4})$ complexity is well established as the standard result for general weakly convex optimization [davis2019stochastic]. The central analytical tool in such results is typically the Moreau envelope, which serves as a potential function in the convergence analysis [davis2019stochastic]. To obtain better complexity guarantees, previous works have leveraged additional structural properties of the objective through smoothing techniques. To the best of our knowledge, all existing smoothing-based approaches for weakly convex functions impose explicit structural assumptions on the objective. For instance, drusvyatskiy2019efficiency studied composite functions of the form $f(\mathbf{x})=h(F(\mathbf{x}))+r(\mathbf{x})$, where $h$ is convex and $F$ is a smooth map. They applied the prox-linear method to this composite setup, solving each subproblem inexactly via accelerated gradient methods and achieving an improved complexity of $\mathcal{O}(1/\varepsilon^{3})$. Similarly, bohm2021variable, peng2023riemannian developed a variable smoothing technique for problems of the form $f(\mathbf{x})=g(A\mathbf{x})$, where $g$ is weakly convex and $A$ is a linear operator. However, their analysis relies on the surjectivity of $A$, an assumption that our framework does not require. Notably, this smoothing methodology arises as a special instance within our broader framework. Our proposed approach is more general, accommodating a wider class of weakly convex objectives without imposing restrictive structural assumptions on either the objective function or the smoothing procedure.
@@
The relaxation of global Lipschitz continuity assumptions in weakly convex optimization has received significant attention in recent years. mai2021stability pioneered this direction by analyzing the stability and convergence of stochastic gradient clipping methods beyond Lipschitz continuity and smoothness assumptions. li2024revisiting extended the classical subgradient method to handle non-Lipschitz convex and weakly convex functions, establishing a complexity of $\mathcal{O}(1/\varepsilon^{2})$ for convex objectives and $\mathcal{O}(1/\varepsilon^{4})$ for weakly convex objectives, without requiring modifications to the algorithm or imposing additional assumptions on the subgradients. In the stochastic setting, gao2024stochastic developed adaptive regularization strategies that maintain the $\mathcal{O}(1/\varepsilon^{4})$ complexity for stochastic weakly convex optimization, allowing the Lipschitz parameter to be either a general function of the iterate norm or estimated locally through random samples. zhu2023unified provided a unified analysis framework of subgradient methods for minimizing composite nonconvex, nonsmooth, and non-Lipschitz functions, establishing convergence guarantees under diminishing step sizes.
@@
The use of smoothing techniques extends far beyond weakly convex functions and has a rich and influential history (see, e.g., [zang1980smoothing, xu2009smooth, chen2004smoothing, chen2012smoothing, lin2022gradient, kume2024variable]). Notably, xu2009smooth introduced a smooth sample average approximation (SAA) framework for nonsmooth stochastic optimization, demonstrating that stationary points of the smoothed SAA converge, in a suitable sense, to stationary points of the original nonsmooth problem, and illustrating their results through several concrete applications. chen2012smoothing established a general smoothing theory grounded in the principle of gradient consistency, which unifies many classical smoothing methods and ensures the asymptotic convergence of gradient-type algorithms to Clarke stationary points under mild assumptions. In contrast to these works, the present paper addresses weakly convex (and potentially non-Lipschitz) composite optimization problems and seeks to develop an explicit complexity theory. More recently, lin2022gradient proposed a gradient-free approach to nonsmooth, nonconvex optimization using zeroth-order randomized smoothing. It is important to note that their analysis pertains to more general Lipschitz continuous objectives and establishes convergence only to $\delta$-Goldstein stationary points, a notion weaker than the stationarity considered in this work.
@@ 2 Preliminaries
We use boldface letters (e.g. $\mathbf{x}$ and $\mathbf{y}$) to represent vectors. Let $\mathbb{R}^{d}$ be a $d$-dimensional Euclidean space with Euclidean inner product $\langle\cdot,\cdot\rangle$; we use $\lVert\mathbf{x}\rVert$ to express the induced norm $\lVert\cdot\rVert=\sqrt{\langle\cdot,\cdot\rangle}$. We use $\lVert\cdot\rVert_{p}$ to denote the vector $L^{p}$-norm. Hence $\lVert\cdot\rVert_{2}$ is $\lVert\cdot\rVert$ by default. For a set $\mathcal{S}\subset\mathbb{R}^{d}$, we define $\mathrm{dist}(\mathbf{x},\mathcal{S})\coloneqq\inf\{\lVert\mathbf{y}-\mathbf{x}\rVert:\mathbf{y}\in\mathcal{S}\}$ and use $\lVert\mathcal{S}\rVert\coloneqq\mathrm{dist}(\mathbf{0},\mathcal{S})$ to denote its distance to the origin. For any convex function $f:\mathbb{R}^{d}\rightarrow(-\infty,+\infty]$, its Fenchel conjugate is defined by $f^{*}(\mathbf{y})\coloneqq\sup_{\mathbf{x}}\{\langle\mathbf{y},\mathbf{x}\rangle-f(\mathbf{x})\}$. When $f$ is closed and convex, we have biconjugacy $f=f^{**}$ [borwein2006convex, 4.2]. Denote the proximal operator of $f$ by $\mathrm{prox}_{f/\beta}(\mathbf{x})\coloneqq\operatornamewithlimits{arg\,min}_{\mathbf{y}\in\mathbb{R}^{d}}\{f(\mathbf{y})+\frac{\beta}{2}\lVert\mathbf{y}-\mathbf{x}\rVert^{2}\}$. We use $\mathrm{lev}(f,v)\coloneqq\{\mathbf{x}\in\mathbb{R}^{d}:f(\mathbf{x})\leq v\}$ to denote the $v$-sublevel set of a function $f$.
@@
Let $f$ be a $\rho$-weakly convex function. The Moreau envelope [moreau1965proximite] of $f$ is defined as the smooth function
@@
Condition (5) strictly generalizes the classical $L$-Lipschitz smoothness assumption: if $\mathcal{L}(\mathbf{x},\mathbf{y})\equiv L$ for some constant $L>0$, $f$ is said to be an $L$-Lipschitz smooth function on $\mathcal{X}$. The requirement that $\mathcal{L}(\mathbf{x},\mathbf{y})$ be symmetric is made without loss of generality; if an initial $\mathcal{L}$ is not symmetric, it can be replaced by $\min\{\mathcal{L}(\mathbf{x},\mathbf{y}),\mathcal{L}(\mathbf{y},\mathbf{x})\}$.
@@ 3 Smoothing theory
This section develops the theory of smooth approximation for weakly convex optimization. We begin by formalizing the notion of a smooth approximation of a function and establishing its connection to the stationarity measures of the original function. As an important application of our framework, we discuss partial smoothing for composite optimization problems.
@@ 2. S2:
Remark 3.1.
Conditions S1-S2 allow both the approximation error $\mathcal{R}_{\eta}(\mathbf{x})$ and the local smoothness modulus $\mathcal{L}_{\eta}(\mathbf{x},\mathbf{y})$ to be local, i.e., dependent on the evaluation point(s). This flexibility is crucial for handling nonsmooth functions that are not globally Lipschitz continuous, thereby broadening the class of problems that can be handled. Section 4.1 elaborates more on this aspect.
The framework above generalizes the smooth approximation theory for convex optimization [Beck2012smoothing], which assumes $f_{\eta}$ to be convex and globally Lipschitz smooth. In the convex setting, convergence rates are typically measured by the function value gap. Condition S1 allows one to directly relate the optimality gap of the nonsmooth problem to that of its smooth approximation. Unfortunately, this connection does not help for nonconvex optimization, as achieving global optimality is generally intractable. Therefore, we instead establish a relationship between the first-order properties of the approximation $f_{\eta}$ and those of the original function $f$. We start by showing that the gradient of the smoothed function can be interpreted as a specific approximate subgradient of $f$, which is defined below.
@@
If $f$ is convex, then a $(0,\varepsilon)$-subgradient reduces to the $\varepsilon$-subgradient in convex optimization [brondsted1965subdifferentiability]. To our knowledge, its extension to weakly convex functions was previously explored by ruszczynski1987linearization, and its use in the complexity analysis for nonconvex optimization has been more recently leveraged in boob2024level and van2024weak. Our definition differs slightly by allowing the weak convexity parameter of the approximation $\bar{\rho}$ to be different from $\rho$. This flexibility allows us to handle smoothing operations that might increase the weak convexity modulus (i.e. $\bar{\rho}>\rho$), a scenario we will examine further in the next section.
@@
Let $f$ be proper, lower semi-continuous, and $\rho$-weakly convex, and let $\mathbf{x}\in\operatornamewithlimits{dom}f$.
@@
$=f(\mathbf{x})+\langle\hat{\mathbf{u}},\mathbf{y}-\mathbf{x}\rangle+\langle\hat{\mathbf{u}}-\mathbf{u},\mathbf{x}-\hat{\mathbf{x}}\rangle-\frac{\rho}{2}\lVert\mathbf{y}-\hat{\mathbf{x}}\rVert^{2}-\frac{\rho}{2}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert^{2}.$ (12)
@@
Placing this result in (12), we have

### v3

@@ 1 Introduction
Here, $f$ is a nonsmooth nonconvex function, and $r$ is a convex, lower-bounded, and lower semi-continuous function. We assume that $f$ is $\rho$-weakly convex for some $\rho>0$; that is, the function $f(\mathbf{x})+\frac{\rho}{2}\|\mathbf{x}\|^{2}$ is convex. We also assume that $r$ is prox-friendly, meaning its associated proximal operator can be computed efficiently. Weakly convex optimization problems frequently arise in various data-driven applications, including signal processing [17], machine learning [8], and reinforcement learning [40].
The nonsmoothness of $f$ prevents the direct application of classical complexity results for smooth nonconvex optimization [20, 21]. Early progress in weakly convex optimization was achieved using double-loop proximal algorithms that require solving auxiliary subproblems at each iteration [14, 16]. A major breakthrough came from Davis and Drusvyatskiy [13], who proved that single-loop methods, such as the (stochastic) subgradient method, achieve an $\mathcal{O}(1/\varepsilon^{4})$ complexity bound for finding an approximate stationary point. Their work reveals an implicit smoothing effect in nonsmooth optimization: algorithms like the proximal subgradient method can be interpreted as optimizing the Moreau envelope $\phi^{\hat{\rho}}(\mathbf{x})$ ($\hat{\rho}>\rho$), a smooth approximation of $\phi$. Notably, $\phi$ and its Moreau envelope $\phi^{\hat{\rho}}$ share the same stationary points, and the gradient of the envelope provides a natural measure of approximate stationarity for (1). This insight has established the Moreau envelope as a central tool in the algorithm analysis for weakly convex problems, influencing a wide range of subsequent work [30, 15, 19, 9, 28, 43].
@@
Explicit smoothing has a long history [11]. In particular, Nesterov’s pioneering work [34] showed that applying accelerated methods to a suitably smoothed approximation of a structured convex problem improves the complexity of finding an $\varepsilon$-optimal solution from $\mathcal{O}(1/\varepsilon^{2})$ to $\mathcal{O}(1/\varepsilon)$. Building on this idea, Beck and Teboulle [2] developed a unified smoothing framework that integrates multiple smoothing schemes with accelerated gradient methods. However, both [34] and [2] restrict their analyses to convex problems. While earlier works have examined the asymptotic behavior of smoothing in nonconvex settings [11, 41], a rigorous complexity theory for explicit smoothing in the weakly convex regime has remained underdeveloped.
@@ 1.1 Contributions
To address this gap, we extend the smooth approximation framework of Beck and Teboulle [2] from the convex setting to the broader class of weakly convex problems. In particular, our framework seamlessly incorporates two prominent smoothing techniques as special cases: 1) Moreau envelope smoothing, and 2) Nesterov-type smoothing for compositions of convex functions with smooth mappings. A key advantage of this generalization arises in stochastic optimization, where one can smooth the function associated with each stochastic sample individually and then perform minibatching. This strategy can provide computational benefits over model-based minibatching algorithms [13, 15], which require solving a potentially expensive proximal subproblem at every iteration. Another important consequence is that Moreau envelope smoothing becomes substantially more practical: it can now be used not only for convergence analysis [13], but also as an effective and scalable algorithmic tool.
It is important to emphasize that the smooth approximation we propose, akin to that of Beck and Teboulle [2], describes a broad class of surrogate functions and is not tied to any particular smoothing technique. We establish quantitative relationships connecting the stationarity measures of the smoothed problem (2) to those of the original nonsmooth problem (1). A key implication is that the design of the smooth surrogate can be decoupled from the design of the optimization algorithm. This stands in contrast to many prior smoothing strategies for weakly convex optimization, which often rely on exploiting structural properties of the underlying problem [4, 16].
@@
A further contribution of our work is the relaxation of the global Lipschitz smoothness assumption [2] in the design of smooth approximation functions. As a result, we can handle objectives that are not globally Lipschitz, thereby substantially broadening the class of problems to which our framework applies. For example, in real phase retrieval, the loss $\ell(\mathbf{x})=|\langle\mathbf{a},\mathbf{x}\rangle^{2}-b|$ has an unbounded subgradient in $\mathbb{R}^{d}$, and standard Nesterov-type smoothing produces approximations that fail to satisfy global Lipschitz smoothness. To overcome this limitation, we employ a line search-based accelerated gradient method for convex optimization problems without global Lipschitz smoothness. The use of line search ensures that accelerated convergence rates for convex problems remain attainable even when global smoothness is absent. Under mild assumptions, we show that the inexact proximal point method, equipped with this accelerated solver for its subproblems, still achieves an $\mathcal{O}(1/\varepsilon^{3})$ complexity guarantee. This result markedly improves upon the previous $\mathcal{O}(1/\varepsilon^{4})$ complexity achieved by non-Lipschitz subgradient methods [23, 43]. Finally, our theoretical findings are supported by numerical experiments. Our results demonstrate that smoothing leads to smoother convergence behavior and outperforms standard subgradient methods on robust nonlinear regression problems.
@@ 1.2 Related works
The $\mathcal{O}(1/\varepsilon^{4})$ complexity is well established as the standard result for general weakly convex optimization [13]. The central analytical tool in such results is typically the Moreau envelope, which serves as a potential function in the convergence analysis [13]. To obtain better complexity guarantees, previous works have leveraged additional structural properties of the objective through smoothing techniques. To the best of our knowledge, all existing smoothing-based approaches for weakly convex functions impose explicit structural assumptions on the objective. For instance, Drusvyatskiy and Paquette [16] studied composite functions of the form $f(\mathbf{x})=h(F(\mathbf{x}))+r(\mathbf{x})$, where $h$ is convex and $F$ is a smooth map. They applied the prox-linear method to this composite setup, solving each subproblem inexactly via accelerated gradient methods and achieving an improved complexity of $\mathcal{O}(1/\varepsilon^{3})$. Similarly, Böhm and Wright [4], Peng et al. [36] developed a variable smoothing technique for problems of the form $f(\mathbf{x})=g(A\mathbf{x})$, where $g$ is weakly convex and $A$ is a linear operator. However, their analysis relies on the surjectivity of $A$, an assumption that our framework does not require. Notably, these previous smoothing techniques arise as special instances within our broader framework. Our proposed approach is more general, accommodating a wider class of weakly convex objectives without imposing restrictive structural assumptions on either the objective function or the smoothing procedure.
@@
The relaxation of global Lipschitz continuity assumptions in weakly convex optimization has received significant attention in recent years. Mai and Johansson [31] pioneered this direction by analyzing the stability and convergence of stochastic gradient clipping methods beyond Lipschitz continuity and smoothness assumptions. Li et al. [28] extended the classical subgradient method to handle non-Lipschitz convex and weakly convex functions, establishing a complexity of $\mathcal{O}(1/\varepsilon^{2})$ for convex objectives and $\mathcal{O}(1/\varepsilon^{4})$ for weakly convex objectives, without requiring modifications to the algorithm or imposing additional assumptions on the subgradients. In the stochastic setting, Gao and Deng [18] developed adaptive regularization strategies that maintain the $\mathcal{O}(1/\varepsilon^{4})$ complexity for stochastic weakly convex optimization, allowing the Lipschitz parameter to be either a general function of the iterate norm or estimated locally through random samples. Zhu et al. [43] provided a unified analysis framework of subgradient methods for minimizing composite nonconvex, nonsmooth, and non-Lipschitz functions, establishing convergence guarantees under diminishing step sizes.
@@
The use of smoothing techniques extends far beyond weakly convex functions and has a rich and influential history (see, e.g., [42, 41, 10, 11, 29, 25]). Notably, Xu and Zhang [41] introduced a smooth sample average approximation (SAA) framework for nonsmooth stochastic optimization, demonstrating that stationary points of the smoothed SAA converge, in a suitable sense, to stationary points of the original nonsmooth problem, and illustrating their results through several concrete applications. Chen [11] established a general smoothing theory based on gradient consistency, which unifies many classical smoothing methods and ensures the asymptotic convergence of gradient-type algorithms to Clarke stationary points under mild assumptions. In contrast to these works, the present paper addresses weakly convex (and potentially non-Lipschitz) composite optimization problems and seeks to develop an explicit complexity theory. More recently, Lin et al. [29] proposed a gradient-free approach to nonsmooth, nonconvex optimization using zeroth-order randomized smoothing. It is important to note that their analysis pertains to more general Lipschitz continuous objectives and establishes convergence only to $\delta$-Goldstein stationary points, a notion weaker than the stationarity considered in this work.
@@ 2 Preliminaries
We use boldface letters (e.g. $\mathbf{x}$ and $\mathbf{y}$) to represent vectors. Let $\mathbb{R}^{d}$ be a $d$-dimensional Euclidean space with Euclidean inner product $\langle\cdot,\cdot\rangle$; we use $\lVert\mathbf{x}\rVert$ to express the induced norm $\lVert\cdot\rVert=\sqrt{\langle\cdot,\cdot\rangle}$. We use $\lVert\cdot\rVert_{p}$ to denote the vector $L^{p}$-norm. Hence $\lVert\cdot\rVert_{2}$ is $\lVert\cdot\rVert$ by default. For a set $\mathcal{S}\subset\mathbb{R}^{d}$, we define $\mathrm{dist}(\mathbf{x},\mathcal{S})\coloneqq\inf\{\lVert\mathbf{y}-\mathbf{x}\rVert:\mathbf{y}\in\mathcal{S}\}$ and use $\lVert\mathcal{S}\rVert\coloneqq\mathrm{dist}(\mathbf{0},\mathcal{S})$ to denote its distance to the origin. For any convex function $f:\mathbb{R}^{d}\rightarrow(-\infty,+\infty]$, its Fenchel conjugate is defined by $f^{*}(\mathbf{y})\coloneqq\sup_{\mathbf{x}}\{\langle\mathbf{y},\mathbf{x}\rangle-f(\mathbf{x})\}$. When $f$ is closed and convex, we have biconjugacy $f=f^{**}$ [6, 4.2]. Denote the proximal operator of $f$ by $\mathrm{prox}_{f/\beta}(\mathbf{x})\coloneqq\operatornamewithlimits{arg\,min}_{\mathbf{y}\in\mathbb{R}^{d}}\{f(\mathbf{y})+\frac{\beta}{2}\lVert\mathbf{y}-\mathbf{x}\rVert^{2}\}$. We use $\mathrm{lev}(f,v)\coloneqq\{\mathbf{x}\in\mathbb{R}^{d}:f(\mathbf{x})\leq v\}$ to denote the $v$-sublevel set of a function $f$.
@@
Let $f$ be a $\rho$-weakly convex function. The Moreau envelope [33] of $f$ is defined as the smooth function
@@
Condition (5) strictly generalizes the classical global $L$-Lipschitz smoothness assumption: if $\mathcal{L}(\mathbf{x},\mathbf{y})\equiv L$ for some constant $L>0$, $f$ is said to be an (globally) $L$-Lipschitz smooth function on $\mathcal{X}$. The requirement that $\mathcal{L}(\mathbf{x},\mathbf{y})$ be symmetric is made without loss of generality; if an initial $\mathcal{L}$ is not symmetric, it can be replaced by $\min\{\mathcal{L}(\mathbf{x},\mathbf{y}),\mathcal{L}(\mathbf{y},\mathbf{x})\}$.
@@ 3 Smoothing theory
This section develops the theory of smooth approximation for weakly convex optimization. We begin by formalizing the notion of smooth approximation of a function and establishing its connection to the stationarity measures of the original function. As an important application of our framework, we discuss partial smoothing for composite optimization problems.
@@ 2. S2:
Conditions S1S2 differ from the standard notion of smooth approximation [2], in which $\mathcal{R}_{\eta}$ and $\mathcal{L}_{\eta}$ do not depend on $\mathbf{x}$ or $\mathbf{y}$. This added flexibility is crucial for handling nonsmooth functions that are not globally Lipschitz continuous, thereby broadening the class of problems that can be addressed; see Section 4.1. Moreover, our framework extends smooth approximation theory beyond the convex setting to encompass nonconvex functions. In the convex setting, convergence rates are typically measured in terms of the function value gap, and Condition S1 directly relates the optimality gap of the nonsmooth problem to that of its smooth approximation. This connection is less informative in nonconvex optimization, where achieving global optimality is generally intractable. Accordingly, we instead establish a relationship between the first-order properties of the approximation $f_{\eta}$ and those of the original function $f$. We start by showing that the gradient of the smoothed function can be interpreted as a specific approximate subgradient of $f$, which is defined below.
@@
If $f$ is convex, then a $(0,\varepsilon)$-subgradient reduces to the $\varepsilon$-subgradient in convex optimization [7]. To our knowledge, its extension to weakly convex functions was previously explored by Ruszczyński [38], and its use in the complexity analysis for nonconvex optimization has been more recently leveraged in Boob et al. [5] and van Ackooij et al. [39]. Our definition differs slightly by allowing the weak convexity parameter of the approximation $\bar{\rho}$ to be different from $\rho$. This flexibility allows us to handle smoothing operations that might increase the weak convexity modulus (i.e. $\bar{\rho}>\rho$), a scenario we will examine further in the next section.
@@
Let $f$ be proper, lower semi-continuous, and $\rho$-weakly convex $\rho>0$, and let $\mathbf{x}\in\operatornamewithlimits{dom}f$.
@@
$=f(\mathbf{x})+\langle\hat{\mathbf{u}},\mathbf{y}-\mathbf{x}\rangle+\langle\hat{\mathbf{u}}-\mathbf{u},\mathbf{x}-\hat{\mathbf{x}}\rangle-\frac{\rho}{2}\lVert\mathbf{y}-\hat{\mathbf{x}}\rVert^{2}-\frac{\rho}{2}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert^{2}.$
@@
Placing this result in (3.1), we have